OpenStudy (anonymous):

Solve for x: 3(x + 4) = 39

OpenStudy (anonymous):

distribute the 3 to the (x+4) so it'd be 3x+12= 39. then subtract 12 and divide by 3

OpenStudy (mimi_x3):

3x + 12 = 39 3x = 39-12 3x = 27 x=9

OpenStudy (anonymous):

x=9